# Concurrent Composition for Differentially Private Continual Mechanisms

### v3

@@ 6.3. Simplified Formulation for Pseudocodes
Decision vs. Auxiliary Variables.
We partition the variables of a continual mechanism $\mathcal{M}$ into two categories:
  • Decision variables: These are either outputs of submechanisms or variables that are computed solely based on decision variables. Intuitively, it follows that they are post-processing of the privatized outputs.
  • Auxiliary variables: These are all remaining variables. The values of these variables may depend on the mechanism’s inputs as well as on both decision and auxiliary variables.
@@
For simplicity, we assume pseudocodes are deterministic as any randomness could be modeled as a submechanism call. For instance, sampling from the uniform distribution could be represented as invoking a 0-DP submechanism that outputs such a sample on a fixed input.
Properties.
@@
    Destination property. A pseudocode satisfies this property if (i) the decision of whether to produce an output, create a new submechanism, or interact with an existing one, (ii) the choice of the new submechanism in the second case, and (iii) the selection of the submechanism to interact with in the third case is made solely based on decision variables.
@@
    Response property. A pseudocode satisfies this property if its output is determined solely by decision variables.
@@
    $f\to f^{\prime}$ property. Let $f$ and $f^{\prime}$ be verification functions. A pseudocode satisfies this property if, for every $t\in\mathbb{N}$ and every pair of input sequences $(\sigma^{0},\sigma^{1})$ of length $t$ forming an $f$-valid combination, the following holds: Assume that all decision variables of $\mathcal{M}$ are updated identically through the end of time step $t$. For each $b\in\{0,1\}$, define $\sigma^{b}$ to denote the sequence of creation queries and inputs to submechanisms when processing $\sigma^{b}$. Here, a creation query is a tuple specifying a created submechanisms, its privacy parameters, and associated verification functions. Also, an input $m$ to the $j$-th submechanism is represented by the pair $(m,j)$. Then, the combination of $\sigma^{0}$ and $\sigma^{1}$ is required to be $f^{\prime}$-valid.
@@ 7. Application (Continual Monotone Histogram Query Problem)
Let $\sim$ be a binary relation on a family of message sequences. Throughout this section, we say that a mechanism is $(\epsilon,\delta)$-DP w.r.t. the neighbor relation $\sim$ (or simply it is $(\epsilon,\delta)$-DP) if the mechanism is $(\epsilon,\delta)$-DP w.r.t. to the verification function $f_{\sim}$ against adaptive adversaries, where $f_{\sim}$ is defined as follows: For every message sequence $m_{1},\dots,m_{t}$, $f_{\sim}(m_{1}\dots,m_{t})=\top$ if and only if the message sequence is suitable (implying $m_{j}=(m_{j}^{0},m_{j}^{1})$ for every $j\in\{1,\dots,t\}$) and the sequences $\sigma_{0}=(m_{1}^{0},\dots,m_{t}^{0})$ and $\sigma_{1}=(m_{1}^{1},\dots,m_{t}^{1})$ satisfy $\sigma_{0}\sim\sigma_{1}$.
@@
Let $d,T\in\mathbb{N}$ and let $q:\mathbb{N}^{d}\to\mathbb{N}$ be a function that is monotone in its input. In the (differentially private) continual monotone histogram query problem, a continual mechanism $\mathcal{M}$ receives $T$ inputs of row vectors $x^{1},\dots,x^{T}$, each in $\{0,1\}^{d}$. Define $\sim$ to be a neighbor relation on $\left(\{0,1\}^{d}\right)^{*}$ where two input streams are neighboring if they are identical except for a single step. Differential privacy for $\mathcal{M}$ is defined with respect to this neighbor relation (see Remark 7.1). The output of $\mathcal{M}$ at each time step $t$ is (an additive approximation to) $q(t)=q((\sum_{\ell=1}^{t}x_{i}^{\ell})_{i\in[d]})$.
@@
Recall function $q:\{0,1\}^{d}\to\mathbb{R}$ and neighbor relation $\sim$ from Definition 7.2. $q$ is said to be $1$-sensitive if $\max_{x^{0}\sim x^{1}}|q(x^{0})-q(x^{1})|$ is at most $1$.
Henzinger, Sricharan, and Steiner [13] presented the best known differentially private mechanism for the monotone histogram query problem, where the query is $1$-sensitive. For $\epsilon$-differential privacy it has, with probability at least $1-\beta$, an additive error of $O((d\log^{2}(dq^{*}/\beta)+\log T)\epsilon^{-1})$, and for $(\epsilon,\delta)$-differential privacy, it has an additive error of $O((d\log^{1.5}(dq^{*}/\beta)+\log T)\epsilon^{-1}\log\frac{1}{\delta})$, where $q^{*}=\max_{t\in[T]}q(t)$. The privacy of the mechanism is proven “from scratch” using the concept of a privacy game [14] and the latter bound only holds against an oblivious adversary.
In Section 7.1, we give an overview of the mechanism in [13], which we will call HSS, and in Section 7.2, we give a simple privacy proof for it using our approach. With the new privacy analysis, the result for $(\epsilon,\delta)$-differential privacy is stronger and holds against an adaptive adversary.
@@
To clarify the connection between these pseudocodes and our definition of continual mechanisms (Definition 3.1), we note that the state of $\textrm{SVT}_{h,\epsilon,q}$ is the tuple of its variables. When $\texttt{SVT-init}(\epsilon,q,h)$ is invoked, the continual mechanism $\textrm{SVT}_{h,\epsilon,q}$ with the single initial state $(\epsilon,q,h,\tau=\text{null})$ is created. In the pseudocode of SVT-init, the variable $\tau$ takes value. In our setting, such initialization is considered to happen when $\textrm{SVT}_{h,\epsilon,q}$ receives its first input. The SVT-update operation then describes how $\textrm{SVT}_{h,\epsilon,q}$ maps its current state (i.e., set of variables) and an input message $m=(x,\texttt{Thresh})\in\{0,1\}^{d}\times\mathbb{R}$ to a new states (with updated variables) and a response $m^{\prime}\in\{\top,\bot\}$.
@@
As part of their privacy proof, the following is shown in [13]:
@@
If $q$ is a $1$-sensitive function, then for every $\epsilon$ and $h$, the mechanism $\textrm{SVT}_{\epsilon/6,q,h}$ is $\epsilon/3$-DP w.r.t. the neighbor relation $\sim$ in Definition 7.2.
$d$-dimensional continual counter. The mechanism d-counter is described by two operations. The first, $\texttt{d-counter-init}(\epsilon)$, creates $d$ differentially private continual binary counters with privacy parameter $\epsilon/d$. The second, $\texttt{d-counter-update}(x)$, gives the elements of its $d$-dimensional input $x$ to their corresponding binary counters and returns the aggregated outputs as a $d$-dimensional vector. We denote this continual mechanism by $\texttt{d-counter}_{\epsilon}$. In another version, $\texttt{d-counter-init}(\epsilon,\delta)$ creates $d$ differentially private continual binary counters with privacy parameters $\epsilon/d$ and $\delta/d$. We represent this continual mechanism by $\texttt{d-counter}_{\epsilon,\delta}$.
@@
Mechanisms $\texttt{d-counter}_{\epsilon}$ and $\texttt{d-counter}_{\epsilon,\delta}$ are $\epsilon$-DP and $(\epsilon,\delta)$-DP, respectively.
Continual histogram query mechanism. The mechanism HSS, described in Algorithm 4, is initialized with a privacy parameter $\epsilon>0$ and an accuracy parameter $0<\beta\leq 1$ and receives a vector $x\in\{0,1\}^{d}$ as input at each time step. This mechanism partitions the input stream into subsequences, called intervals. At the end of each interval, it forwards the $d$-dimensional sum of all the inputs of the interval to the differentially private $d$-dimensional continual counter d-counter. (Note that d-counter can be used directly to solve the problem; however, HSS achieves better accuracy bounds in many cases.) The mechanism HSS maintains two $d$-dimensional vectors: (1) variable $c$, where $c_{i}$ stores the exact sum of coordinate $i$ within the current interval, and (2) variable $h$, where $h_{i}$ stores the approximate sum of coordinate $i$ over all previous intervals computed by d-counter. Both vectors are initialized to $0^{d}$.
@@
The mechanism HSS also stores a threshold variable Thresh, initially set to $\gamma(1,1,\beta,\epsilon)$. Upon receiving an input $x$ at step $t$, the mechanism updates $c$ and invokes $\texttt{SVT-update}(x,\texttt{Thresh})$. If the response is $\top$, the current interval ends: $c$ is passed to d-counter, $h$ is updated with the output of d-counter, and $c$ is reset to $0^{d}$. Since SVT halts after returning $\top$, a new instance $\textrm{SVT}_{\epsilon/6,q,h}$ is created. In this case, HSS privately evaluates $q(c+h)$ via the Laplace noise and compares the noisy result to $\texttt{Thresh}-\xi(t,j,\beta,\epsilon)$. If the inequality holds, Thresh is suitably increased. Moreover, independent of the outcome of this comparison, Thresh is updated again to reflect the increase in the number of intervals. Then, regardless of the output of $\texttt{SVT-update}(x,\texttt{Thresh})$, HSS updates Thresh once more to account for the increased number of time steps. Finally, HSS outputs $h$.
@@
  $c_{i}\leftarrow c_{i}+x_{i}$ for all $i\in[d]$,
  if $\texttt{SVT-update}(x,\texttt{Thresh})==\top$ then
@@
     $\texttt{Thresh}=\texttt{Thresh}+\gamma(t,j,\beta,\epsilon)$
   end if
@@
In Algorithm 4, the summation $q(c+h)+\mathrm{Lap}(3/\epsilon)$ can be rewritten as an instance of Laplace mechanism with privacy parameter $\epsilon/3$ and input $q(c+h)$. Consequently, the mechanism HSS executes:
@@
By Corollary 7.5, d-counter is $\epsilon/3$-DP (or $(\epsilon/3,\delta)$-DP when approximate counters are used). By Lemma 7.4, each SVT instance is $\epsilon/3$-DP. Finally, the Laplace mechanism with parameter $\epsilon/3$ is a non-interactive mechanism (NIM) that is well known to satisfy $\epsilon/3$-DP on neighboring real inputs differing by at most $1$. The idea is to show that the concurrent composition of multiple instances of these mechanisms created by HSS is $\epsilon$-DP (or $(\epsilon,\delta)$-DP when d-counter is approx DP). Then, we will then show that HSS satisfies the required properties in Corollary 6.5, and thus it has the same privacy guarantees as the concurrent composition. As concurrent compositions theorems in this paper require the composed mechanisms to have discrete answer distributions, we first show that we can assume the output space of d-counter and Laplace mechanism is discrete. Since SVT has a finite output space, its answer distributions are also discrete.
@@
Discretization of outputs. As argued earlier, the counters executed by d-counter outputs integer values, and thus d-counter has discrete answer distributions. Likewise, we can assume that the Laplace mechanism in HSS rounds its noisy real-valued output to the nearest integer. By the post-processing lemma, this rounding does not degrade privacy, and by definition, it increases the additive error by at most $1$, which is unimportant for the accuracy guarantees of [13]. Hence, both d-counter and the Laplace mechanism can be treated as having discrete output distributions.
@@
Pseudocode privacy analysis. Recall from Section 6.3 that decision variables of a mechanism are either (i) outputs of submechanisms, or (ii) variables computed solely as functions of other decision variables. In Algorithm 4, the variables $h$, outout, $j$, $t$, and Thresh are decision variables, while $c$ is not: Variable $h$ denotes the privatized outcomes of d-counter. Variable outout equals $q(h)$, where $h$ is a decision variable. Index $j$ is initially $1$ and is increased by $1$ when the privatized output of SVT is $\top$. Time step $t$ is initially 0 and is increase by $1$ at each step, independent of the input and other variables. Finally, variable Thresh is initialized and updated to values determined by $\epsilon,\beta,j,t$, where $\epsilon,\beta$ are fixed and $j,t$ are decision variables. It is important to note that the decision of updating variables $h$ (and consequently outout) and Thresh depends only on the privatized output of SVT.
@@
We first show that the destination property holds. For that we have to show that the decision of whether to produce an output, create a new submechanism (and if so, which one) or interact with an existing one (and if so, which one) are solely based on decision variables. All submechanisms of HSS are instantiated either in the initial step (and, thus, whether it receives input is independent of non-decision variables and the inputs $x$) or based on outputs of the private submechanism SVT. The active SVT instance receives input at every step, and thus whether it receives input is independent of non-decision variables and the inputs $x$. Also, whether the mechanism d-counter gets updated depends only on the privatized output of SVT. Finally, HSS returns an output after the final update of Thresh, without any condition on the inputs $x$ or the non-decision variables. Thus, the destination property in Corollary 6.5 holds. Moreover, the final output of HSS is outout, which is a decision variable, so the response property in that corollary also holds. It remains to show HSS satisfies the $f\to f^{\prime}$ property, where $f$ is the verification function corresponding to the neighbor relation $\sim$ in Definition 7.2 and $f^{\prime}$ is a verification satisfying that the $f^{\prime}$-concurrent composition of the continual mechanisms is $(\epsilon,\delta)$-DP.
Fix some values for all decision variables at every time step. Conditioned on these values, when HSS processes two neighboring input sequences, the resulting inputs to its submechanisms behave as follows:
  • d-counter receives neighboring sequences,
  • all but one instance of SVT receive identical inputs, with at most one receiving neighboring inputs, and
  • all but one Laplace mechanism receive identical inputs, with at most one receiving neighboring inputs.

### v4

@@ 6.3. Simplified Formulation for Pseudocodes
Tainted vs. Untainted Variables.
To rigorously track sensitive data, we rely on concepts from computer security. Information flow characterizes how data transfers between variables during a program’s execution (denning1976lattice, ; sabelfeld2003language, ). Taint analysis builds upon this to monitor the propagation of sensitive inputs, known as taint sources (newsome2005dynamic, ; schwartz2010all, ). Any variable computed using a taint source becomes a tainted variable. To safely release data, it must be processed by a sanitizer—in our context, a differentially private submechanism. The sanitizer mathematically interrupts the sensitive information flow, rendering its outputs and all subsequent downstream computations as untainted variables.
To formalize this, we partition the variables of a mechanism into two categories using a recursive definition based on their inputs:
  • Tainted variables: The mechanism’s inputs are strictly tainted. Furthermore, any intermediate variable computed via a function where at least one input argument is a tainted variable is also considered tainted. These variables retain a direct dependence on the raw inputs.
  • Untainted variables: Variables computed solely from privatized data or fixed values (such as the privacy parameters or constants) are untainted. Crucially, the output of any differentially private submechanism (the sanitizer) is defined as untainted, regardless of whether its inputs were tainted. Any subsequent variable computed as a function of only untainted variables remains untainted. Consequently, untainted variables are purely a post-processing of privatized outputs.
@@
For simplicity, we assume the pseudocode of $\mathcal{M}$ is deterministic as any randomness could be modeled as a submechanism call. For instance, sampling from the uniform distribution could be represented as invoking a 0-DP submechanism that outputs such a sample on a fixed input.
Definition 6.5.
@@
    Destination property. A pseudocode satisfies this property if (i) the decision of whether to produce an output, create a new submechanism, or interact with an existing one, (ii) the choice of the new submechanism in the second case, and (iii) the selection of the submechanism to interact with in the third case is made solely based on untainted variables.
@@
    Response property. A pseudocode satisfies this property if its output is determined solely by untainted variables.
@@
    $f\to f^{\prime}$ property. Let $f$ and $f^{\prime}$ be verification functions. A pseudocode satisfies this property if, for every $t\in\mathbb{N}$ and every pair of input sequences $(\sigma^{0},\sigma^{1})$ of length $t$ forming an $f$-valid combination, the following holds: Assume that all untainted variables of $\mathcal{M}$ are updated identically through the end of time step $t$. For each $b\in\{0,1\}$, define $\sigma^{b}$ to denote the sequence of creation queries and inputs to submechanisms when processing $\sigma^{b}$. Here, a creation query is a tuple specifying a created submechanism, its privacy parameters, and associated verification functions. Also, an input $m$ to the $j$-th submechanism is represented by the pair $(m,j)$. Then, the combination of $\sigma^{0}$ and $\sigma^{1}$ is required to be $f^{\prime}$-valid.
@@ 7. Application (Continual Monotone Histogram Query Problem)
Let $\sim$ be a binary relation on a family of message sequences. Throughout this section, we say that a mechanism is $(\epsilon,\delta)$-DP w.r.t. the neighbor relation $\sim$ (or simply it is $(\epsilon,\delta)$-DP) if the mechanism is $(\epsilon,\delta)$-DP w.r.t. the verification function $f_{\sim}$ against adaptive adversaries, where $f_{\sim}$ is defined as follows: For every message sequence $m_{1},\dots,m_{t}$, $f_{\sim}(m_{1}\dots,m_{t})=\top$ if and only if the message sequence is suitable in the sense of Definition 4.5 (implying $m_{j}=(m_{j}^{0},m_{j}^{1})$ for every $j\in\{1,\dots,t\}$) and the sequences $\sigma_{0}=(m_{1}^{0},\dots,m_{t}^{0})$ and $\sigma_{1}=(m_{1}^{1},\dots,m_{t}^{1})$ satisfy $\sigma_{0}\sim\sigma_{1}$.
@@
Let $d,T\in\mathbb{N}$ and let $q:\mathbb{N}^{d}\to\mathbb{N}$ be a function that is monotone in its input. In the (differentially private) continual monotone histogram query problem, a continual mechanism $\mathcal{M}$ receives $T$ inputs of row vectors $x=(x^{1},\dots,x^{T})$, each in $\{0,1\}^{d}$. Differential privacy for $\mathcal{M}$ is defined with respect to the neighbor relation $\sim_{H}$ defined as follows (see Remark 7.1). For $T\in\mathbb{N}$, two input streams in $\left(\{0,1\}^{d}\right)^{T}$ are neighboring if they are identical except for a single step. The output of $\mathcal{M}$ at each time step $t$ is (an additive approximation to) $q(t)=q((\sum_{\ell=1}^{t}x_{i}^{\ell})_{i\in[d]})$.
@@
Recall function $q:\{0,1\}^{d}\to\mathbb{R}$ and neighbor relation $\sim_{H}$ from Definition 7.2. $q$ is said to be $1$-sensitive if $\max_{x^{0}\sim_{H}x^{1}}|q(x^{0})-q(x^{1})|$ is at most $1$.
Henzinger, Sricharan, and Steiner (henzingerSS2023, ) presented the best known differentially private mechanism for the monotone histogram query problem, where the query is $1$-sensitive. For $\epsilon$-differential privacy it has, with probability at least $1-\beta$, an additive error of $O((d\log^{2}(dq^{*}/\beta)+\log T)/\epsilon)$, and for $(\epsilon,\delta)$-differential privacy, it has an additive error of $O((d\log^{1.5}(dq^{*}/\beta)+\log T)\log\frac{1}{\delta}/\epsilon)$, where $q^{*}=\max_{t\in[T]}q(t)$. The privacy of the mechanism is proven “from scratch” using the concept of a privacy game (pricedpco, ) and the bound was only shown for $(\epsilon,\delta)$-DP against an oblivious adversary.
In Section 7.1, we give an overview of the mechanism in (henzingerSS2023, ), which we will call HSS, and in Section 7.2, we give a simple privacy proof for it using our approach. With the new privacy analysis, the result for $(\epsilon,\delta)$-differential privacy is stronger and holds against an adaptive adversary.
@@
To clarify the connection between these pseudocodes and our definition of continual mechanisms (Definition 3.1), we note that the state of $\textrm{SVT}_{h,\epsilon,q}$ is the tuple of its variables. When $\texttt{SVT-init}(\epsilon,q,h)$ is invoked, the continual mechanism $\textrm{SVT}_{h,\epsilon,q}$ with the single initial state $(\epsilon,q,h,\tau=\text{null})$ is created. In the pseudocode of SVT-init, the variable $\tau$ takes a value. In our setting, such initialization is considered to happen when $\textrm{SVT}_{h,\epsilon,q}$ receives its first input. The SVT-update operation then describes how $\textrm{SVT}_{h,\epsilon,q}$ maps its current state (i.e., set of variables) and an input message $m=(x,\texttt{Thresh})\in\{0,1\}^{d}\times\mathbb{R}$ to a new state (with updated variables) and a response $m^{\prime}\in\{\top,\bot\}$.
@@
As part of their privacy proof, the following is shown in (henzingerSS2023, ):
@@
If $q$ is a $1$-sensitive function, then for every $\epsilon$ and $h$, the mechanism $\textrm{SVT}_{\epsilon/6,q,h}$ is $\epsilon/3$-DP w.r.t. the neighbor relation $\sim_{S}$ defined as follows (see Remark 7.1). For $T\in\mathbb{N}$, two input streams $\left((x_{1},\tau_{1}),\dots,(x_{T},\tau_{T})\right)$ and $\left((y_{1},\nu_{1}),\dots,(y_{T},\nu_{T})\right)$ in $\left((\{0,1\}^{d}\times\mathbb{R})\right)^{T}$ are neighboring if the (threshold) sequences $(\tau_{1},\dots,\tau_{T})$ and $(\nu_{1},\dots,\nu_{T})$ are identical at every entry and the (binary) sequences $(x_{1},\dots,x_{T})$ and $(y_{1},\dots,y_{T})$ are identical in all but one entry.
$d$-dimensional continual counter. The mechanism d-counter is described by two operations. The first, $\texttt{d-counter-init}(\epsilon)$, creates $d$ continual counters that are $\epsilon/d$-DP w.r.t. the neighbor relation $\sim^{\prime}$, where two real-valued sequences of the same length are neighbors if they differ only on a single step by at most $1$. The second, $\texttt{d-counter-update}(c)$, gives the elements of its $d$-dimensional input $c\in\mathbb{R}^{d}$ to their corresponding counters and returns the aggregated outputs as a $d$-dimensional vector. We denote this continual mechanism by $\texttt{d-counter}_{\epsilon}$. In another version, $\texttt{d-counter-init}(\epsilon,\delta)$ creates $d$ continual counters, each satisfying $(\epsilon/d,\delta/d)$-DP w.r.t. $\sim^{\prime}$. We represent this continual mechanism by $\texttt{d-counter}_{\epsilon,\delta}$.
@@
Mechanisms $\texttt{d-counter}_{\epsilon}$ and $\texttt{d-counter}_{\epsilon,\delta}$ are $\epsilon$-DP and $(\epsilon,\delta)$-DP w.r.t. the neighbor relation $\sim_{C}$ defined as follows (see Remark 7.1). For $T\in\mathbb{N}$, two input streams $x=(x_{1},\dots,x_{T})$ and $y=(y_{1},\dots,y_{T})$ in $(\mathbb{R}^{d})^{T}$ are neighboring if there exists $i\in[T]$ such that $x_{j}=y_{j}$ for all $j\in[T]\setminus\{i\}$ and $|x_{i}-y_{i}|_{\infty}\leq 1$, i.e., each entry of $x_{i}$ and $y_{i}$ differs by at most $1$.
Continual histogram query mechanism. The mechanism HSS, described in Algorithm 4, is initialized with a privacy parameter $\epsilon>0$, a query $q:\mathbb{N}^{d}\to\mathbb{N}$, and an accuracy parameter $0<\beta\leq 1$ and receives a vector $x\in\{0,1\}^{d}$ as input at each time step. This mechanism partitions the input stream into subsequences, called intervals. At the end of each interval, it forwards the $d$-dimensional sum of all the inputs of the interval to the differentially private $d$-dimensional continual counter d-counter. (Note that d-counter can be used directly to solve the problem; however, HSS achieves better accuracy bounds in many cases.) The mechanism HSS maintains two $d$-dimensional vectors: (1) variable $c$, where $c_{i}$ stores the exact sum of coordinate $i$ within the current interval, and (2) variable $h$, where $h_{i}$ stores the approximate sum of coordinate $i$ over all previous intervals computed by d-counter. Both vectors are initialized to $0^{d}$.
@@
The mechanism HSS also stores a threshold variable Thresh, initially set to $\gamma(1,1,\beta,\epsilon)$. Upon receiving an input $x$ at step $t$, the mechanism updates $c$ and invokes $\texttt{SVT-update}(x,\texttt{Thresh})$. If the response, denoted by $a$, is $\top$, the current interval ends and HSS takes the following actions: (threshold update) It first privately evaluates $q(c+h)$ via the Laplace noise and compares the noisy result to $\texttt{Thresh}-\xi(t,j,\beta,\epsilon)$. If the inequality holds, Thresh is suitably increased. Then independent of the outcome of this comparison, Thresh is updated again to reflect the increase in the number of intervals. (summation update) Subsequently, HSS passes $c$ to d-counter, updates $h$ with the output of d-counter, and resets $c$ to $0^{d}$. Since SVT halts after returning $\top$, HSS also creates a new instance $\textrm{SVT}_{\epsilon/6,q,h}$.
Finally, regardless of the output of $\texttt{SVT-update}(x,\texttt{Thresh})$, HSS updates Thresh once more to account for the increased number of time steps and outputs $h$.
@@
  $c_{i}\leftarrow c_{i}+x_{i}$ for all $i\in[d]$
  $a=\texttt{SVT-update}(x,\texttt{Thresh})$
  if $a==\top$ then
     $\texttt{Thresh}=\texttt{Thresh}+\gamma(t,j,\beta,\epsilon)$
   end if
@@
In Algorithm 4, the summation $q(c+h)+\mathrm{Lap}(3/\epsilon)$ can be rewritten as an instance of Laplace mechanism with privacy parameter $\epsilon/3$ and input $q(c+h)$. The privacy of the Laplace mechanism is defined with respect to the neighbor relation $\sim_{L}$, where two real values $r_{1},r_{2}\in\mathbb{R}$ satisfy $r_{1}\sim_{L}r_{2}$ if and only if $|r_{1}-r_{2}|\leq 1$. Consequently, the mechanism HSS executes:
@@
By Corollary 7.5, d-counter is $\epsilon/3$-DP (or $(\epsilon/3,\delta)$-DP when approximate counters are used) w.r.t. $\sim_{C}$. By Lemma 7.4, each SVT instance is $\epsilon/3$-DP w.r.t. $\sim_{S}$. Finally, the Laplace mechanism with parameter $\epsilon/3$ is a non-interactive mechanism (NIM) that is well known to satisfy $\epsilon/3$-DP on neighboring real inputs differing by at most $1$ (i.e., w.r.t. to neighbor relation $\sim_{L}$ defined above). The idea is to show that the concurrent composition of multiple instances of these mechanisms created by HSS is $\epsilon$-DP (or $(\epsilon,\delta)$-DP when d-counter is approx DP). Then, we will show that HSS satisfies the required properties in Corollary 6.6, and thus it has the same privacy guarantees as the concurrent composition. As concurrent composition theorems in this paper require the composed mechanisms to have discrete answer distributions, we first show that we can assume the output space of d-counter and Laplace mechanism is discrete. Since SVT has a finite output space, its answer distributions are also discrete.
@@
Discretization of outputs. As argued earlier, the counters executed by d-counter output integer values, and thus d-counter has discrete answer distributions. Likewise, we can assume that the Laplace mechanism in HSS rounds its noisy real-valued output to the nearest integer. By the post-processing lemma, this rounding does not degrade privacy, and by definition, it increases the additive error by at most $1$, which is unimportant for the accuracy guarantees of (henzingerSS2023, ). Hence, both d-counter and the Laplace mechanism can be treated as having discrete output distributions.
@@
Pseudocode privacy analysis. Recall from Section 6.3 that untainted variables of a mechanism are (i) outputs of submechanisms, (ii) fixed values (e.g., privacy and accuracy parameters), or (iii) variables computed solely as functions of other untainted variables. In Algorithm 4, the variables $\epsilon$, $\beta$, $q$, $a$, $h$, outout, $j$, $t$, and Thresh are untainted, while $c$ is not: The values of the variables $\epsilon$, $\beta$, and $q$ are fixed and independent of the sensitive data. Variables $a$ and $h$ denote the privatized outcomes of SVT and d-counter, respectively. Variable outout equals $q(h)$, where $h$ is an untainted variable. Index $j$ is initially $1$ and is increased by $1$ when $a=\top$. Time step $t$ is initially 0 and is increased by $1$ at each step, independent of the input and other variables. Finally, variable Thresh is initialized and updated to values determined by the untainted variables $\epsilon,\beta,j,t$. It is important to note that the decision of updating variables $h$ (and consequently outout) and Thresh depends only on the privatized output of SVT, i.e., the untainted variable $a$.
@@
We first show that the destination property holds. For that we have to show that the decision of whether to produce an output, create a new submechanism (and if so, which one) or interact with an existing one (and if so, which one) is solely based on untainted variables. All submechanisms of HSS are instantiated either in the initial step (and, thus, whether it receives input is independent of tainted variables and the inputs $x$) or based on outputs of the private submechanism SVT. The active SVT instance receives input at every step, and thus whether it receives input is independent of tainted variables and the inputs $x$. Also, whether the mechanism d-counter gets updated depends only on the privatized output of SVT. Finally, HSS returns an output after the final update of Thresh, without any condition on the inputs $x$ or the tainted variables. Thus, the destination property in Corollary 6.6 holds. Moreover, the final output of HSS is outout, which is an untainted variable, so the response property in that corollary also holds. It remains to show HSS satisfies the $f\to f^{\prime}$ property, where $f$ is the verification function corresponding to the neighbor relation $\sim_{H}$ in Definition 7.2 and $f^{\prime}$ is a verification function satisfying that the $f^{\prime}$-concurrent composition of the continual mechanisms is $(\epsilon,\delta)$-DP.
Let $T\in\mathbb{N}$, and let $x=(x_{1},\dots,x_{T})$ and $y=(y_{1},\dots,y_{T})$ be two sequences in $\left(\{0,1\}^{d}\right)^{T}$ satisfying $x\sim_{H}y$, i.e., there exists $i\in[T]$ such that $x_{j}=y_{j}$ for all $j\in[T]\setminus\{i\}$. Fix arbitrary values for all untainted variables at every time step. Conditioned on these values, we compare the input streams of HSS’s submechanisms when this mechanisms runs on $x$ and $y$:
  • Inputs of d-counter: By construction, the input stream of HSS is partitioned into intervals determined by the time steps at which SVT outputs $\top$. Since these outputs are untainted, they are identical in both executions to the fixed values specified in the condition. Let $1\leq t_{1}<\dots<t_{k}\leq T$ be the steps SVT outputs $\top$, and set $t_{0}=0$. The submechanism d-counter receives the interval sums $\left(\sum_{\ell=t_{j-1}+1}^{t_{j}}x_{\ell}\right)_{j=1}^{k}$ and $\left(\sum_{\ell=t_{j-1}+1}^{t_{j}}y_{\ell}\right)_{j=1}^{k}$ as input sequence when HSS runs on $x$ and $y$, respectively. All intervals not containing $i$ have identical sums. If $i>t_{k}$, no (closed) interval includes $i$ and thus all sums coincide. If $i\leq t_{k}$, let $j^{*}\in[k]$ denote the index satisfying $t_{j^{*}-1}+1\leq i\leq t_{j^{*}}$. Then
    $\sum_{\ell=t_{j^{*}-1}+1}^{t_{j^{*}}}x_{\ell}-\sum_{\ell=t_{j^{*}-1}+1}^{t_{j^{*}}}y_{\ell}=x_{i}-y_{i}.$
    Since $x_{i},y_{i}\in\{0,1\}^{d}$, we have $x_{i}-y_{i}\in\{-1,0,1\}^{d}$ and thus $\|x_{i}-y_{i}\|_{\infty}\leq 1$. Hence the inputs to d-counter differ in at most one coordinate by at most one, i.e., they are $\sim_{C}$-neighbors (see Corollary 7.5).
  • Inputs of SVT: The threshold variable Thresh is untainted. Let $\texttt{Thresh}_{\ell}$ denote its fixed value at time $\ell$. Let time steps $t_{1},\dots,t_{k}$ be defined as above. Each time SVT outputs $\top$, the mechanism HSS reinitializes it. Therefore, HSS initiates $k+1$ instances of SVT. Let $t_{0}=0$ and $t_{k+1}=T$. For each $j\in[k+1]$, the $j$-th instance of SVT receives $\left((x_{\ell},\texttt{Thresh}_{\ell})\right)_{\ell=t_{j-1}+1}^{t_{j}}$ and $\left((y_{\ell},\texttt{Thresh}_{\ell})\right)_{\ell=t_{j-1}+1}^{t_{j}}$ as input sequence when HSS runs on $x$ and $y$, respectively. All instances except possibly $j^{*}$ (the one containing $i$) receive identical inputs. Let $j^{*}\in[k+1]$ denote the index where $t_{j^{*}-1}+1\leq i\leq t_{j^{*}}$. For every $j\in[k+1]\setminus\{j^{*}\}$, the input sequences of the $j$-th SVT submechanism are identical when HSS runs on $x$ and $y$. Moreover, the input sequences of the $j^{*}$-th SVT submechanism are identical except for the inputs $(x_{i},\texttt{Thresh}_{i})$ and $(y_{i},\texttt{Thresh}_{i})$, which have the same threshold $\texttt{Thresh}_{i}$. Hence, these input sequences are $\sim_{S}$-neighbors (see Lemma 7.4).
  • Inputs of the Laplace Mechanism: The Laplace mechanism is applied to the value $q(h+c)$, where $c$ is the interval sum. The variable $h$ is untainted and therefore fixed and identical in both executions of HSS. As shown above, all interval sums coincide except possibly one, and that one differs by at most $1$ in $\ell_{\infty}$-norm. Hence the corresponding values $h+c$ also differ by at most $1$ in $\ell_{\infty}$-norm. Since $q$ is $1$-sensitive (see Definition 7.3), the real-valued outputs of $q$ on these inputs differ by at most $1$. Therefore, when HSS runs on $x$ and $y$, all corresponding Laplace submechanisms receive identical inputs, except possibly one pair whose inputs differ by at most $1$, that is, they are $\sim_{L}$-neighbors.